Weiland Co. shows the following information on its 2016 income statement: sales = $160,000; costs = $80,500; other expenses = $3,800; depreciation expense = $9,500; interest expense = $7,000; taxes = $20,720; dividends = $7,900. In addition, you're told that the firm issued $4,000 in new equity during 2016 and redeemed $6,700 in outstanding long-term debt.
  
a. What is the 2016 operating cash flow? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
  
Operating cash flow            $
  
b. What is the 2016 cash flow to creditors? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
  
Cash flow to creditors            $
  
c. What is the 2016 cash flow to stockholders? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
  
Cash flow to stockholders            $
  
d. If net fixed assets increased by $20,950 during the year, what was the addition to NWC? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
  
Addition to net working capital    
Answer #1

Net income=( sales-cost of goods sold-depreciaiton-other-interest-taxes)
=(160000-80500-9500-3800-7000-20720)
=38480
a)operating cash flow=EBIT+depreciation-taxes
EBIT=sales-cost of goods sold-depreciaiton-other exp
=(160000-80500-9500-3800
=66200
operating cash flow=66200+9500-20720=54980
b)Cash flow to creditors = Interest – net borrowing
=7000-(-6700)
=13700
c)Cash flow to stockholder= Dividends paid-new equity issues
=7900-4000
=3900
d)Cash flow from assets = OCF – Change in Net working capital – Net capital spending
Change in Net working capital=OCF –Net capital spendingl-cash flow from assets
Cash flow from assets = Cash flow to creditors + Cash flow to stockholders
=13700+3900=17600
Net capital spending = Depreciation + Increase in fixed assets
=9500+20950=30450
Change in Net working capital=54980-30450-17600=6930